# Index and Robustness of Mixed Equilibria: an algebraic approach

### 04 Mar 2026 v1

@@ Abstract.
I am grateful to Srihari Govindan and Rida Laraki for comments and stimulating conversations.
@@ 1. Introduction
One of the cleanest criteria for equilibrium selection stemming out of this literature is the non-zero index rule. [^1] The index is an integer number associated to an equilibrium that captures the robustness of the equilibrium to perturbations of its associated best-reply correspondence to nearby continuous maps. Whenever the number is non-zero, the equilibrium is robust in this sense. Non-zero index equilibria are in particular robust to payoff perturbations, and are, for example, proper. Non-zero index equilibria have recently been axiomatized, for generic-games in extensive form, by Govindan and Wilson (2025) [^2] The axiomatization is of decision-theoretic nature and allows for a direct comparison to the solution concept of Mertens stability, also axiomatized in the same paper., and Pahl and Pimienta (2025) have proved a simple characterization that shows that these equilibria are hyperstable (Kohlberg and Mertens (1986)), i.e., they are the invariant and payoff-robust ones. [^3] Invariance here means that if we add finitely many duplicate strategies to the game, that is, mixed strategies as pure strategies, the hyperstable equilibria are still payoff-robust in the game with added strategies. Payoff-robustness of an equilibrium means that for sufficiently small payoff-perturbations, there are equilibria close to the given equilibrium point.
The topological tools involved in the computation of the index of equilibria do not fully explore the fact that the equilibria of finite games are described by polynomial systems of equalities and inequalities. In fact, when considering the standard formulae for computing the index of equilibria (see subsection 3.6 of Pahl (2023)), the particular polynomial nature of the maps is ignored. There are many difficulties that arise from this. An example: a known method to prove that a certain equilibrium is payoff robust is the verification of its non-zero index (this is a sufficient condition for payoff-robustness): to verify that, one can perturb the payoffs (in normal form) of the game to a nearby generic payoff function. Only a finite number of equilibria (with index +1 or -1) arise around the original equilibrium, and the index of this equilibrium is the sum of the indices of the equilibria that arose from the perturbation. There are a number of difficulties with the application of this procedure, for instance: how small should the generic payoff-perturbation be? Is there a canonical way to perturb the game in examples? After perturbing the game, is there a generally applicable method to find all equilibria nearby?
@@ 2.2. Topological Preliminaries
Notationwise, the index of a fixed point $x$ of a map $f$ will be denoted ind(x,f)(x,f). The index of an equilibrium $\sigma^{*}$ refers to the index of the isolated equilibrium $\sigma^{*}$ of a game $G$ with respect to any Nash-map that has $\sigma^{*}$ as fixed point when the game $G$ is fixed (see subsection 2.2.2 in Pahl (2023) for details on Nash-maps). For an open set $U\subset\mathbb{R}^{m}$ such that the closure $\bar{U}$ contains a unique solution $x$ of the equation $p(x)=0\in\mathbb{R}^{m}$, we denote by $\text{deg}_{0}(x,p)$ the topological degree of x w.r.t. to $p$. When the map $p$ is understood, we omit reference to $p$. Recall that the degree and the index of an isolated equilibrium are identical (the result actually holds for connected components of equilibria). For two players, the result is proved in Govindan and Wilson (1997). The general result is proved in DeMichelis and Germano (2000).
@@
$|\text{deg}_{0}(f)|=\text{dim}_{\mathbb{R}}A-2\text{dim}_{\mathbb{R}}I_{max},$
@@ 3. Index and Robustness
Each class of $\mathbb{R}[[u,w]]/\langle u,h(u,w)\rangle$ can be represented by a power series in $w$ only, since each variable in $u$ is zeroed by the ideal $\langle u,h(u,w)\rangle$. This implies that we have an $\mathbb{R}$-algebra isomorphism $\mathbb{R}[[u,w]]/\langle h(0,w)\rangle\cong\mathbb{R}[[u,w]]/\langle u,h(u,w)\rangle$. Therefore we have obtained that $\mathbb{R}[[X_{1},....,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle\cong\mathbb{R}[[w]]/\langle g\rangle$, where $g=h(0,\cdot)$ as $\mathbb{R}$-algebras. This concludes the proof of the“only if” part of the Lemma 3.5.
@@
In order to apply Theorem 5 in Datta (2003) and obtain a finite game associated to the zero-set of each of the polynomials above, we change variables so that the unique root 0 is translated to the point where all coordinates are equal to $1/3$, satisfying the requirement of that theorem. Of course, the topological degree of 0 in either case is invariant under this translation. Applying Theorem 5, we obtain the required game. Now, one needs to verify that the topological degrees of the equilibria are maintained under the new system describing the completely mixed equilibria. For that, consider $F$ as one the polynomial maps constructed above. We checked that in the case of $F(x,y)=(x^{2},y)$, the algebra $\mathbb{R}[[X,Y]]/\langle X^{2},Y\rangle$ is finite dimensional, and a similar argument yields the same finite-dimensionality for $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ in the case $F$ is specified as the other polynomials. Applying Theorem 5 in Datta, this implies that the absolute value of the topological degree of the equilibria is maintained, but one needs to show that the signs are maintained as well, so we need a stronger result. We use Theorem 1.2 in Eisenbud et al. (1977), which provides a formula for the topological degree based on the signature of the bilinear form defined over the $\mathbb{R}$-algebra. Proposition 2 in Eisenbud (1978) yields that the minimal non-zero ideal in $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ is generated by the class of the Jacobian determinant of this polynomial map. Since $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ is isomorphic to $\mathbb{R}[[X_{1},...,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle$ (the $\mathbb{R}$-algebra associated to the system of completely mixed equilibria associated to $F$ and obtained by Theorem 5 in Datta), that minimal non-zero ideal maps to the minimal non-zero ideal of $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ under the isomorphism. This implies that if $\phi:\mathbb{R}[[X_{1},...,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle\to\mathbb{R}$ is a linear functional that takes value $1$ in the class of the Jacobian determinant, then the signature of the bilinear form $a,b]:=\phi(ab),a,b\in\mathbb{R}[[X_{1},...,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle$ is equal to the topological degree of 0 under $p$. The isomorphism between the algebras, call it $\varphi$, allows us to define the linear functional $\psi=\phi\circ\varphi^{-1}$ on $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ that takes value $1$ on the class of the Jacobian determinant of $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$. Note that the signaure of $\cdot$$\cdot$] equals the signature of $(\cdot,\cdot)$, by construction. Hence, by Theorem 1.2 in Eisenbud and Levine (ibid.), the signature of $(\cdot,\cdot)$ is equal to the topological degree of 0 under $F$. The proof is then concluded by observing that the topological degree and index of the equilibrium point agree, by Proposition 3.1. ∎

### 17 Mar 2026 v2

@@ Abstract.
I am grateful to Srihari Govindan and Rida Laraki for comments and stimulating conversations. I thank Julien Fixary for comments and proofreading.
@@ 1. Introduction
One of the cleanest criteria for equilibrium selection stemming out of this literature is the non-zero index rule. [^1] The index is an integer number associated to an equilibrium that captures the robustness of the equilibrium to perturbations of its associated best-reply correspondence to nearby continuous maps. Whenever the number is non-zero, the equilibrium is robust in this sense. Non-zero index equilibria are in particular robust to payoff perturbations, and are, for example, proper. Non-zero index equilibria have recently been axiomatized, for generic-games in extensive form, by Govindan and Wilson (2025) [^2] The axiomatization is of decision-theoretic nature and allows for a direct comparison to the solution concept of Mertens stability, also axiomatized in the same paper., and Pahl and Pimienta (2025) have proved a simple characterization that shows that these equilibria are hyperstable (Kohlberg and Mertens (1986)), i.e., they are the invariant and payoff-robust ones. [^3] Invariance here means that if we add finitely many duplicate strategies to the game, that is, mixed strategies as pure strategies, the hyperstable equilibria are still payoff-robust in the game with added strategies. Payoff-robustness of an equilibrium means that for sufficiently small payoff-perturbations, there are equilibria close to the given equilibrium point. In particular, in terms of desirable properties, non-zero index equilibrium outcomes (i.e., isolated equilibrium outcomes induced by a non-zero index component of mixed strategies) satisfy all the properties highlighted in Kohlberg and Mertens (1986): existence, admissibility, invariance, iterated dominance and sequential rationality. Because of these desirable properties, which are not simultaneously satisfied by other classical refinement concepts (perfect, proper, sequential equilibria, etc.), there is a strong case for the selection and computation of non-zero index equilibrium outcomes in finite games.
Unfortunately, in most non-trivial cases, the computation of the index is not an easy task to undertake. The topological tools involved in the computation of the index of equilibria do not fully explore the fact that the equilibria of finite games are described by polynomial systems of equalities and inequalities. In fact, when considering the standard formulae for computing the index of equilibria (see subsection 3.6 of Pahl (2023)), the particular polynomial nature of the maps is ignored. There are many difficulties that arise from this. An example: a known method to prove that a certain equilibrium is payoff robust is the verification of its non-zero index (this is a sufficient condition for payoff-robustness): to verify that, one can perturb the payoffs (in normal form) of the game to a nearby generic payoff function. Only a finite number of equilibria (with index +1 or -1) arise around the original equilibrium, and the index of this equilibrium is the sum of the indices of the equilibria that arose from the perturbation. There are a number of difficulties with the application of this procedure, for instance: how small should the generic payoff-perturbation be? Is there a canonical way to perturb the game in examples? After perturbing the game, is there a generally applicable method to find all equilibria nearby?
@@ 2.2. Topological Preliminaries
Notationwise, the index of a fixed point $x$ of a map $f$ will be denoted ind(x,f)(x,f). The index of an equilibrium $\sigma^{*}$ refers to the index of the isolated equilibrium $\sigma^{*}$ of a game $G$ with respect to any Nash-map that has $\sigma^{*}$ as fixed point when the game $G$ is fixed (see subsection 2.2.2 in Pahl (2023) for details on Nash-maps). For an open set $U\subset\mathbb{R}^{m}$ such that the closure $\bar{U}$ contains a unique solution $x$ of the equation $p(x)=0\in\mathbb{R}^{m}$, we denote by $\text{deg}_{0}(x,p)$ the topological degree of x w.r.t. to $p$. When the point $x$ is understood, we omit reference to it and write $\text{deg}_{0}(p)$. Recall that the degree and the index of an isolated equilibrium are identical (the result actually holds for connected components of equilibria). For two players, the result is proved in Govindan and Wilson (1997). The general result is proved in DeMichelis and Germano (2000).
@@
$|\text{deg}_{0}(0,f)|=\text{dim}_{\mathbb{R}}(\mathbb{R}[[X_{1},....,X_{n}]]/I)-2\text{dim}_{\mathbb{R}}I_{max},$
@@ 3. Index and Robustness
Each class of $\mathbb{R}[[u,w]]/\langle u,h(u,w)\rangle$ can be represented by a power series in $w$ only, since each variable in $u$ is zeroed by the ideal $\langle u,h(u,w)\rangle$. This implies that we have an $\mathbb{R}$-algebra isomorphism $\mathbb{R}[[u,w]]/\langle h(0,w)\rangle\cong\mathbb{R}[[u,w]]/\langle u,h(u,w)\rangle$. Therefore we have obtained that
$\mathbb{R}[[X_{1},....,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle\cong\mathbb{R}[[w]]/\langle g\rangle,$
where $g=h(0,\cdot)$ as $\mathbb{R}$-algebras. This concludes the proof of the“only if” part of the Lemma 3.5.
@@
In order to apply Theorem 5 in Datta (2003) and obtain a finite game associated to the zero-set of each of the polynomials above, we change variables so that the unique root 0 is translated to the point where all coordinates are equal to $1/3$, satisfying the requirement of that theorem. Of course, the topological degree of 0 in either case is invariant under this translation. Applying Theorem 5, we obtain the required game. Now, one needs to verify that the topological degrees of the equilibria are maintained under the new system describing the completely mixed equilibria. For that, consider $F$ as one the polynomial maps constructed above. We checked that in the case of $F(x,y)=(x^{2},y)$, the algebra $\mathbb{R}[[X,Y]]/\langle X^{2},Y\rangle$ is finite dimensional, and a similar argument yields the same finite-dimensionality for $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ in the case $F$ is specified as the other polynomials. Applying Theorem 5 in Datta, this implies that the absolute value of the topological degree of the equilibria is maintained, but one needs to show that the signs are maintained as well, so we need a stronger result. We use Theorem 1.2 in Eisenbud et al. (1977), which provides a formula for the topological degree based on the signature of the bilinear form defined over the $\mathbb{R}$-algebra. Proposition 2 in Eisenbud (1978) yields that the minimal non-zero ideal in $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ is one-dimensional and is generated by the class of the Jacobian determinant of this polynomial map. Since $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ is isomorphic to $\mathbb{R}[[X_{1},...,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle$ (the $\mathbb{R}$-algebra associated to the system of completely mixed equilibria associated to $F$ and obtained by Theorem 5 in Datta), that minimal non-zero ideal maps to the minimal non-zero ideal of $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ under the isomorphism. Call this isomorphism $\varphi$. Since the minimal non-zero ideals of $\mathbb{R}[[X_{1},...,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle$ and $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ are one-dimensional, $\varphi$ maps one to the other, up to a constant $c$. In the particular case of the substitutions implied by Datta’s argument in the proof of Theorem 5, we can take $c=1$. This implies that if $\phi:\mathbb{R}[[X_{1},...,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle\to\mathbb{R}$ is a linear functional that takes value $1$ in the class of the Jacobian determinant, then the signature of the bilinear form $a,b]:=\phi(ab),a,b\in\mathbb{R}[[X_{1},...,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle$ is equal to the topological degree of 0 under $p$. The isomorphism $\varphi$ allows us to define the linear functional $\psi=\phi\circ\varphi^{-1}$ on $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ that takes value $1$ on the class of the Jacobian determinant of $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$. Note that the signaure of $\cdot$$\cdot$] equals the signature of $(\cdot,\cdot)$, by construction. By Theorem 1.2 in Eisenbud and Levine (ibid.), the signature of $(\cdot,\cdot)$ is equal to the topological degree of 0 under $F$. The proof is then concluded by observing that the topological degree and index of the equilibrium point agree, by Proposition 3.1. ∎